OpenStudy (anonymous):

Please Help i been working all day and playing catch up @ganeshie8 if anyone else could help that would be awesome

OpenStudy (anonymous):

Solve each of the following problems using a system of equations. Show all of your work and state your solution in a complete sentence. Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? (2 points) Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? (2 points) Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (2 points) Follow the 5 steps below to complete this problem. (4 points) Step 1: Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following: Drive a boat Drive a jet ski Step 2: Select a current speed of the water in mph. Step 3: Select the number of hours (be reasonable please) that your friend or family member drove the boat or jets ski against the current speed you chose in step 2. Step 4: Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling with the current). Step 5: Write out the word problem you created and calculate how fast your friend or family member was traveling in still water. Round your answer to the nearest mph.

OpenStudy (anonymous):

1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? (2 points) SOLUTION: Let tony age = t years, Bel age = 56 - t years 56 - t = 2t + 8 56 - 8 = 2t + t ==> 3t = 48 ==> t = 16 years THUS Tony = 16 years, and Belinda = 40 years <=====> ANSWER 2.Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? (2 points) SOLUTION: Let adults = a, and Students = 150 - a NOW, 45a + 28.5(150 - a) = 4770 45a + 4275 - 28.5a = 4770 ==> 16.5a = 4770 - 4275 = 495 a = 495 / 16.5 = 30, THUS, Adult = 30, STUDENTS = 120 <=============> ANSWER 3.Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (2 points) LET, Dimes = d, and Quarter = 46 - d d/10 + (46 - d) /4 = 7 Multiply both sides by 20, we get, 2d + 5(46 - d) = 140 ==> 2d + 230 - 5d = 140 -3d = 140 - 230 = -90 ==> d = 30 THUS, Dimes = 30, and Quarters = 16 <==============> ANSWER

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

thanks heather

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

you look like Marilyn Monroe btw :O

OpenStudy (anonymous):

that is Marilyn Monroe lol

OpenStudy (anonymous):

ohh you put a picture of her i thought that was you hahaha

OpenStudy (anonymous):

im a koala jus sayin

OpenStudy (anonymous):

ya I had a pic of me up but their are some creeps on here

OpenStudy (anonymous):

yea their are O.O

ganeshie8 (ganeshie8):

i never did anything... just for the record :|

OpenStudy (anonymous):

Nope ur all good ganeshie :)

OpenStudy (anonymous):

huh? im not a smart Koala :(

OpenStudy (anonymous):

what did she mean

OpenStudy (anonymous):

I ment he wasn't a creep to me

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

o.O LOL

OpenStudy (anonymous):

i love this website

OpenStudy (anonymous):

me to its very helpful and their cool people to tlak to too

OpenStudy (anonymous):

i know right!!

OpenStudy (anonymous):

sayin haha

OpenStudy (anonymous):

well now im gunna do my history so i have free time to play candy crush haha

OpenStudy (anonymous):

lmao if you need help you can tag me :)

OpenStudy (anonymous):

kk thanks :)

OpenStudy (anonymous):

No problem :)